# Joint State and Noise Covariance Estimation

### v1

@@ Abstract
This paper tackles the problem of jointly estimating the noise covariance matrix alongside primary parameters (such as poses and points) from measurements corrupted by Gaussian noise. In such settings, the noise covariance matrix determines the weights assigned to individual measurements in the least squares problem. We show that the joint problem exhibits a convex structure and provide a full characterization of the optimal noise covariance estimate (with analytical solutions) within joint maximum a posteriori and likelihood frameworks and several variants. Leveraging this theoretical result, we propose two novel algorithms that jointly estimate the primary parameters and the noise covariance matrix. To validate our approach, we conduct extensive experiments across diverse scenarios and offer practical insights into their application in robotics and computer vision estimation problems with a particular focus on SLAM.
@@ I Introduction
Due to these challenges, many applications rely on ad hoc noise covariances, such as arbitrary isotropic (or diagonal) covariances, which are either manually set by experts or determined through trial-and-error tuning. Despite being recognized as one of the most critical and widely acknowledged challenges in SLAM [12, Sections III.B, III.G, and V], the problem of noise covariance estimation remains unsolved and understudied in robotics and computer vision literature.
This paper addresses the problem of jointly estimating the (primary) parameters and noise covariance matrices from noisy measurements in a general setting that covers a wide range of estimation problems in robotics and computer vision. For the first time, our algorithms enable automatic estimation of the noise covariance matrix online (i.e., during deployment) without the need for a separate calibration stage with ground truth information. Furthermore, our MAP framework allows the users to incorporate prior information about the noise covariance, such as manufacturer-provided noise calibration.
Our main contributions are as follows:
  1. 1.
    Formulate the joint MAP and ML estimation problems along with several variants that incorporate different structural assumptions on the true noise covariance matrix. To the best of our knowledge, this is the first time such an approach has been proposed in the SLAM literature.
  2. 2.
    Provide analytical expressions for (conditionally) optimal noise covariances (under MAP and ML criteria, and their constrained variants) in Theorem 1.
  3. 3.
    Building on our theoretical results, we propose two (local) types of optimization methods to solve the joint estimation problem and present a rigorous convergence analysis (Theorems 2 and 3). Our algorithms can be easily incorporated into popular (sparse) nonlinear least squares solvers such as [2, 10, 17] as we demonstrate in our experimental results.
  4. 4.
    Offer practical insights for noise covariance estimation in sparse graph-structured problems such as SLAM, and present extensive experimental and ablation studies for linear(ized)-Gaussian problems and pose-graph optimization (PGO).
@@
We use $[n]$ to denote the set of integers from $1$ to $n$. The abbreviated notation $x_{1:n}$ is used to denote $x_{1},\ldots,x_{n}$. The zero matrix (and vector) is denote by 00 where the size should be clear from the context. $\mathbb{S}^{d}_{\succeq 0}$ and $\mathbb{S}^{d}_{\succ 0}$ denote the sets of $d\times d$ symmetric positive semidefinite and positive definite real matrices, respectively. For two symmetric real matrices $A$ and $B$, $A\succeq B$ (resp., $A\succ B$) means $A-B$ is positive semidefinite (resp., positive definite). $A_{ij}$ denotes the $(i,j)$ element of matrix $A$, and $\mathrm{Diag}(A)$ denotes the diagonal matrix obtained by zeroing out the off-diagonal elements of $A$. The standard (Frobenius) inner product between $n\times n$ real matrices $A$ and $B$ is denoted by $\langle A,B\rangle\triangleq\mathrm{trace}(A^{\top}B)$. The Frobenius norm of $A$ is denoted by $\|A\|=\sqrt{\langle A,A\rangle}$. The weighted Euclidean norm of $x$ given a weight matrix $W\succ 0$ is denoted by $\|x\|_{W}\triangleq\sqrt{x^{\top}Wx}$. The probability density function of the multivariate normal distribution of random variable $x$ with mean vector $\mu$ and covariance matrix $\Sigma$ is denoted by $\mathcal{N}(x;\mu,\Sigma)$.
@@ II Related Works
We refer the reader to [6, 11, 12, 24] for comprehensive reviews of state-of-the-art estimation frameworks in robotics and computer vision. Sparse nonlinear least squares solvers for solving these estimation problems can be found in [2, 10, 17]. However, all of these works assume that the noise covariance is known beforehand. In contrast, our approach simultaneously estimates both the primary parameters (e.g., robot’s trajectory) and the noise covariance matrix directly from noisy measurements. The importance of automatic hyperparameter tuning has recently gained recognition in the SLAM literature; see, e.g., [12, Section V] and [13, 14]. We share this perspective and present, to the best of our knowledge, the first principled approach for optimal noise covariance estimation in SLAM and related problems.
@@ II-A Optimal Covariance Estimation via Convex Optimization
ML estimation of the mean and covariance from independent and identically distributed (i.i.d.) Gaussian samples using the sample mean and sample covariance is a classic example found in textbooks. Boyd and Vandenberghe [5, Chapter 7.1.1] show that covariance estimation in this standard setting and several of its variants can be formulated as convex optimization problems. However, many estimation problems that arise in robotics and other engineering disciplines extend beyond the standard setting. While noise samples are assumed to be i.i.d. for each measurement type (Section IV-C1), the measurements themselves are not identically distributed. Each measurement follows a Gaussian distribution with the corresponding noise covariance matrix and a unique mean that depends on an unknown parameter belonging to a manifold. Furthermore, the measurement function varies across different measurements and is often nonlinear. We demonstrate that the noise covariance estimation problem in this more general setting can also be formulated as a convex optimization problem. Similar to [5], we explore several problem variants that incorporate prior information and additional structural constraints on the noise covariance matrix. These variants differ from those studied in [5, Chapter 7.1.1] and admit analytical (closed-form) solutions.
II-B Iterated Generalized Least Squares
[…]
@@ II-C Noise Covariance Estimation in Kalman Filtering
The problem of identifying process and measurement noise models in Kalman filtering has been extensively studied since 1970s; see, e.g., [27, 9, 15] and references therein. The Kalman filter (KF) is equivalent to the recursive MAP estimator in linear-Gaussian models. Similarly, nonlinear variants of the KF, such as the extended Kalman filter (EKF), can be viewed as approximate MAP estimators. Despite conceptual similarities between problems and proposed solutions, Kalman filtering methods are specifically tailored for linear(ized) models in filtering estimation problems within a recursive framework. Consequently, they are not directly applicable to the smoothing problems (or batch settings) addressed in this paper, which represent the state-of-the-art approach in SLAM and related computer vision problems; see, e.g., [11].
Figure 1: Confidence ellipses for 10 samples drawn from the Wishart distribution $\mathcal{W}(P;V,\nu)$ with different parameters. The scale matrix $V$ is set to identity in the left and middle plots, and to a correlated matrix in the right plot. The degrees of freedom $\nu$ are set to $2$ (left) and $10$ (middle and right). As $\nu$ increases, the samples become more concentrated. Increasing $\nu$ changes the scale of samples as well (cf. left and middle).
@@ III Problem Statement
In this paper, we refer to $x_{\text{true}}$ as the primary “parameters” to distinguish them from $\Sigma_{\text{true}}$. To simplify the discussion, we first consider the case where all measurements share the same noise distribution, meaning there is a single noise covariance matrix $\Sigma_{\text{true}}$ in (1). See Section IV-C1 for extensions to more general cases. In robotics and computer vision applications, $\mathcal{M}$ is typically a (smooth) product manifold comprised of $\mathbb{R}^{d}$ and $\operatorname{SO}(d)$ components ($d\in\{2,3\}$) and other real components (e.g., time offsets, IMU biases). We assume the measurement functions $h_{i}:\mathcal{M}\to\mathbb{R}^{m}$ are smooth. This standard model (along with extensions in Section IV) is quite general, capturing many estimation problems such such as SLAM (with various sensing modalities and variants), PGO, point cloud registration (with known correspondences), perspective-$n$-point, and bundle adjustment.
In this paper, we are interested in the setting where the noise covariance matrix $\Sigma_{\text{true}}\succ 0$ is unknown and must be estimated jointly with $x_{\text{true}}$ based on the collected measurements $\{z_{i}\}_{i=1}^{k}$. For convenience, we formulate the problem in the information form and estimate the noise information (or precision) matrix $P_{\text{true}}\triangleq\Sigma_{\text{true}}^{-1}$. Without loss of generality, we assume a non-informative prior on $x_{\text{true}}$ which is almost always the case in real-world applications (effectively treating it as an unknown “parameter”). We assume a Wishart prior on the noise information matrix $P_{\text{true}}$ and denote its probability density function with $\mathcal{W}(P;V,\nu)$ where $V\in\mathbb{S}^{m}_{\succ 0}$ is the scale matrix and the integer $\nu\geq m+1$ is the number of degrees of freedom; Figure 1 illustrates random samples drawn from $\mathcal{W}(P;V,\nu)$. The Wishart distribution is the standard choice for prior in Bayesian statistics for estimating the information matrix from multivariate Gaussian data (in part due to conjugacy); see, e.g., [3, Eq. (2.155)]. In Algorithm 1, we propose a procedure for setting the parameters of the prior, $V$ and $\nu$, when the user has access to a prior estimate $\Sigma_{0}$ for the noise covariance matrix $\Sigma_{\text{true}}$ (e.g., from prior calibration by the manufacturer of the sensor); see Appendix A for a justification.
@@
6:  $V\leftarrow(w_{\text{prior}}\,k\,\Sigma_{0})^{-1}$
7:  $\nu\leftarrow w_{\text{prior}}\,k+m+1$
@@
$&\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}&&\,\,F(x,P)\\ &\text{subject to}&&\text{$P$ is diagonal}.\end{aligned}$ (6)
In the second variant, we constrain the eigenvalues of $\Sigma=P^{-1}$ to $[\lambda_{\text{min}},\lambda_{\text{max}}]$ where $\lambda_{\text{max}}\geq\lambda_{\text{min}}>0$. These eigenvalues specify the minimum and maximum variance along all directions in $\mathbb{R}^{m}$ (i.e., variance of all normalized linear combinations of noise components), Therefore, this constraint allows the user to incorporate prior knowledge about the sensor’s noise limits. We will also show that in many real-world instances (especially with a weak or no prior $w_{\text{prior}}\to 0$), constraining the smallest eigenvalue of $\Sigma$ (or largest eigenvalue of $P$) is essential for preventing $\Sigma$ from collapsing to zero.
@@
$&\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}&&\,\,F(x,P)\\ &\text{subject to}&&\lambda_{\text{max}}^{-1}I\preceq P\preceq\lambda_{\text{min}}^{-1}I.\end{aligned}$ (7)
@@
$&\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}&&\,\,F(x,P)\\ &\text{subject to}&&\lambda_{\text{max}}^{-1}I\preceq P\preceq\lambda_{\text{min}}^{-1}I,\\ &&&\text{$P$ is diagonal}.\end{aligned}$ (8)
Remark 1.
It is easy to verify that by fixing $P$ to a predetermined value, Problem 1 reduces to the standard (weighted) nonlinear least squares over $x$ and can be (locally) solved using existing solvers [2, 10, 17]. [^2] Note that $\langle S(x),P\rangle=\sum_{i=1}^{k}\|r_{i}(x)\|^{2}_{P}$. Similarly, algebraically the problem reduces to joint ML estimation when $w_{\text{prior}}\to 0$ where $w_{\text{prior}}$ is the prior weight in Algorithm 1.333The prior $w_{\text{prior}}$ cannot be exactly zero due to matrix inversion in line 6 of Algorithm 1. To be more precise, $w_{\text{prior}}=0$ refers to the case where $V^{-1}$ is set to zero in $M(x)$ (5) and $\nu=m+1$. This abuse of language allows us to save space by treating ML estimation as a special of the MAP problems introduced in Problems 1-4. In that case, the linear term in the objective becomes $\langle S(x),P\rangle$. See also [26] for the unconstrained ML formulation.
Remark 2.
Singularity of $S(x)$ may lead to an ill-conditioned problems when the prior is weak (i.e., $w_{\text{prior}}$ is small) as the corresponding objective function becomes very sensitive in certain directions that correspond to the smallest eigenvalues of $M(x)$. In the unconstrained joint ML estimation case (i.e., $w_{\text{prior}}=0$ and without $\Sigma\succeq\lambda_{\text{min}}I$), the minimization problem is ill-posed (unbounded from below) when $S(x)$ is singular. [^4] This can be shown by inspecting the objective value along $P+c\cdot u_{0}u_{0}^{\top}$ where $u_{0}$ is in the nullspace of $S(x)$ and $c$ is an arbitrarily large positive scalar. See Remark 5 for more information regarding scenarios where $S(x)$ may become singular.

### v3

@@ Abstract
This paper tackles the problem of jointly estimating the noise covariance matrix alongside states (parameters such as poses and points) from measurements corrupted by Gaussian noise and, if available, prior information. In such settings, the noise covariance matrix determines the weights assigned to individual measurements in the least squares problem. We show that the joint problem exhibits a convex structure and provide a full characterization of the optimal noise covariance estimate (with analytical solutions) within joint maximum a posteriori and likelihood frameworks and several variants. Leveraging this theoretical result, we propose two novel algorithms that jointly estimate the primary parameters and the noise covariance matrix. Our BCD algorithm can be easily integrated into existing nonlinear least squares solvers, with negligible per-iteration computational overhead. To validate our approach, we conduct extensive experiments across diverse scenarios and offer practical insights into their application in robotics and computer vision estimation problems with a particular focus on SLAM.
@@ I Introduction
Due to these challenges, many applications rely on ad hoc noise covariances, such as arbitrary isotropic (or diagonal) covariances, which are either manually set by experts or determined through trial-and-error tuning. Despite being recognized as one of the most critical and widely acknowledged challenges in SLAM [15, Sections III.B, III.G, and V], the problem of noise covariance estimation remains unsolved and understudied in robotics and computer vision literature.
We present, to the best of our knowledge, the first algorithms for online (i.e., during deployment) joint ML/MAP estimation of states and noise covariance matrices from noisy measurements and, when available, prior information. Our approach is general and eliminates the need for a separate calibration stage across a broad class of estimation problems beyond SLAM. We analyze the convergence properties of the proposed algorithm and demonstrate that it can be seamlessly integrated into existing sparse nonlinear least squares solvers [3, 13, 20], with negligible computational overhead.
@@
We use $[n]$ to denote the set of integers from $1$ to $n$. The abbreviated notation $x_{1:n}$ is used to denote $x_{1},\ldots,x_{n}$. The zero matrix (and vector) is denote by 0 where the size should be clear from the context. $\mathbb{S}^{d}_{\succeq 0}$ and $\mathbb{S}^{d}_{\succ 0}$ denote the sets of $d\times d$ symmetric positive semidefinite and positive definite real matrices, respectively. For two symmetric real matrices $A$ and $B$, $A\succeq B$ (resp., $A\succ B$) means ABA-B is positive semidefinite (resp., positive definite). $A_{ij}$ denotes the $(i,j)$ element of matrix $A$, and $\mathrm{Diag}(A)$ denotes the diagonal matrix obtained by zeroing out the off-diagonal elements of $A$. The standard (Frobenius) inner product between $n\times n$ real matrices $A$ and $B$ is denoted by $\langle A,B\rangle\triangleq\mathrm{trace}(A^{\top}B)$. The Frobenius norm of $A$ is denoted by $\|A\|=\sqrt{\langle A,A\rangle}$. The weighted Euclidean norm of $x$ given a weight matrix $W\succ 0$ is denoted by $\|x\|_{W}\triangleq\sqrt{x^{\top}Wx}$. The probability density function of the multivariate normal distribution of random variable $x$ with mean vector $\mu$ and covariance matrix $\Sigma$ is denoted by $\mathcal{N}(x;\mu,\Sigma)$.
@@ II Related Works
We refer the reader to [9, 14, 15, 30] for comprehensive reviews of state-of-the-art estimation frameworks in robotics and computer vision. Sparse nonlinear least squares solvers for solving these estimation problems can be found in [3, 13, 20]. However, all of these works assume that the noise covariance is known beforehand. In contrast, our approach simultaneously estimates both the primary parameters (e.g., robot’s trajectory) and the noise covariance matrix directly from noisy measurements. The importance of automatic hyperparameter tuning has recently gained recognition in the SLAM literature; see, e.g., [15, Section V] and [16, 17]. We share this perspective and present, to the best of our knowledge, the first principled approach for ML/MAP measurement covariance estimation in SLAM and related problems.
@@ II-A Optimal Covariance Estimation via Convex Optimization
ML estimation of the mean and covariance from independent and identically distributed (i.i.d.) Gaussian samples using the sample mean and sample covariance is a classic example found in textbooks. Boyd and Vandenberghe [8, Chapter 7.1.1] show that covariance estimation in this standard setting and several of its variants can be formulated as convex optimization problems. However, many estimation problems that arise in robotics and other engineering disciplines extend beyond the standard setting. While noise samples are assumed to be i.i.d. for each measurement type (Section IV-C1), the measurements themselves are not identically distributed. Each measurement follows a Gaussian distribution with the corresponding noise covariance matrix and a unique mean that depends on an unknown parameter belonging to a manifold. Furthermore, the measurement function varies across different measurements and is often nonlinear. We demonstrate that the noise covariance estimation problem in this more general setting can also be formulated as a convex optimization problem. Similar to [8], we explore several problem variants that incorporate prior information and additional structural constraints on the noise covariance matrix. These variants differ from those studied in [8, Chapter 7.1.1] and admit analytical (closed-form) solutions.
II-B Covariance Estimation in Robotics and Computer Vision
Zhan et al. [33] propose a joint pose and noise covariance estimation method for the perspective-$n$-point (PnnP) problem in computer vision. Their approach is based on the iterated (or iterative) generalized least squares (IGLS) method (see [28, Chapter 12.5] and references therein), alternating between pose and noise covariance estimation. They report improvements in estimation accuracy ranging from $2\%$ to $34\%$ compared to baseline methods that assume a fixed isotropic noise covariance. To the best of our knowledge, before [33], IGLS had not been applied in robotics or computer vision. [^1] We were unable to find the original reference for IGLS. However, the concept was already known and analyzed in the econometrics literature in 1970s [22]. IGLS is closely related to the feasible generalized least squares (FGLS) method which also has a long history in econometrics and regression. Our work (developed concurrently with [33]) generalizes and extends both IGLS and [33] in several keys ways. First, we prove that the joint ML estimation problem is ill-posed when the sample covariance matrix is singular. This critical problem arises frequently in real-world applications, where the sample covariance matrix can be singular or poorly conditioned (Remark 4). We address this critical issue by (i) constraining the minimum eigenvalue of the noise covariance matrix, and, in our MAP formulation, (ii) imposing a Wishart prior on the noise information matrix. We derive analytical optimal solutions for the noise covariance matrix, conditioned on fixed values of the primary parameters, for MAP and ML joint estimation problems and several of their constrained variants (Theorem 1). These enable the end user to leverage prior information about the noise covariance matrix (from, e.g., manufacturer’s calibration) in addition to noisy measurements to solve the joint estimation problem. We propose several algorithms for solving these joint estimation problems and present a rigorous theoretical analysis of their convergence properties. Our formulation is more general than IGLS and we show how our framework can be extended to heteroscedastic measurements, nonlinear (with respect to) noise models, and manifold-valued parameters. Finally, we provide insights into the application of our approach to “graph-structured” estimation problems [19], such as PGO and other SLAM variants. These problems present additional challenges compared to the PnnP problem due to the increasing number of primary parameters (thousands of poses in PGO vs. a single pose in PnnP) and the sparsity of measurements, which reduce the effective signal-to-noise ratio.
Barfoot et al. [5] and Wong et al. [32] propose an EM-type method for learning the noise covariance matrix as part of a variational inference framework. Similar to our work, they used an Inverse-Wishart prior on the covariance matrix. However, we estimate the covariance matrix by solving the joint MAP/ML estimation problems and their constrained variants. While the covariance estimation formulation and proposed techniques share similarities, our method focuses on widely used MAP/ML point estimation rather than obtaining an analytical approximation of the entire posterior, yielding significantly faster solutions (by up to several orders of magnitude based on the statistics reported in [5]). As a result, unlike [5, 32], our method can be readily integrated into existing nonlinear least squares solvers such as [3, 13, 20] in both online and offline settings with a negligible computational overhead.
Barfoot [4] proposes an algorithm for joint state and covariance MAP estimation (without our structural constraints) in which each measurement is assigned a unique covariance. In [4, Chapter 5.5.3], the conditionally optimal covariances are derived and eliminated analytically from the optimization problem. The resulting reduced objective resembles the Cauchy robust cost, which is minimized via iteratively reweighted least squares. Our Elimination algorithm (Algorithm 2) extends this approach to the more general setting where arbitrary subsets of measurements can share a covariance matrix. In this case, the reduced objective includes log-determinant terms and no longer corresponds to the Cauchy cost. We therefore directly optimize the reduced objective using the limited-memory BFGS method. We also propose a second algorithm, Block Coordinate Descent (Algorithm 3), which is preferable to Elimination as it offers faster convergence, exploits problem sparsity, and leverages existing sparse nonlinear least squares solvers.
Lu et al. [21] introduce a covariance autotuning method for object measurements in SLAM, employing an alternating optimization scheme over states and the variances of a diagonal covariance matrix. Despite similarities between [21] and our BCD algorithm, the cost function in [21] differs from standard MAP/ML formulation. Additionally, unlike [21], our approach does not assume that the covariance matrix is diagonal.
Qadri et al. [27] propose a bilevel optimization framework to learn measurement covariances from a calibration dataset with known ground truth. Specifically, they seek the covariance estimate (outer problem) that minimizes the state estimation (inner problem) error. Unlike [27], our method does not directly minimize the estimation error and thus does not require access to the ground truth. Instead, we jointly estimate both states and covariances directly from observed data (and, optionally, prior information on the covariance) in a joint MAP/ML framework. As a result, our method can be used during deployment to learn measurement covariances based on the collected measurements.
@@ II-C Noise Covariance Estimation in Kalman Filtering
The problem of identifying process and measurement noise models in Kalman filtering (often referred to as adaptive Kalman filtering) has been extensively studied since the late 1960s; see, e.g., [1, 23, 34, 12, 18] and references therein. While our work shares certain similarities with these approaches and their underlying principles, such methods are specifically designed for (approximate, when models are nonlinear) recursive MAP estimation in linear(ized) models within the filtering setting. As a result, they are not readily applicable to batch and smoothing formulations, nonlinear measurement models, sparse large-scale problems, or (nonlinear) manifold-valued states. These features are essential for addressing many estimation problems in robotics and computer vision (see, e.g., state-of-the-art estimation frameworks for SLAM [14]).
Figure 1: Confidence ellipses for 10 samples drawn from the Wishart distribution $\mathcal{W}(P;V,\nu)$ with different parameters. The scale matrix $V$ is set to identity in the left and middle plots, and to a correlated matrix in the right plot. The degrees of freedom $\nu$ are set to 22 (left) and 1010 (middle and right). As $\nu$ increases, the samples become more concentrated. Increasing $\nu$ changes the scale of samples as well (cf. left and middle).
@@ III Problem Statement
In this paper, we refer to $x_{\text{true}}$ as the primary “parameters” to distinguish them from $\Sigma_{\text{true}}$. To simplify the discussion, we first consider the case where all measurements share the same noise distribution, meaning there is a single noise covariance matrix $\Sigma_{\text{true}}$ in (1). See Section IV-C1 for extensions to more general cases. In robotics and computer vision applications, $\mathcal{M}$ is typically a (smooth) product manifold comprised of $\mathbb{R}^{d}$ and $\operatorname{SO}(d)$ components ($d\in\{2,3\}$) and other real components (e.g., time offsets, IMU biases). We assume the measurement functions $h_{i}:\mathcal{M}\to\mathbb{R}^{m}$ are smooth. This standard model (along with extensions in Section IV) is quite general, capturing many estimation problems such as SLAM (with various sensing modalities and variants), PGO, point cloud registration (with known correspondences), perspective-$n$-point, and bundle adjustment.
In this paper, we are interested in the setting where the noise covariance matrix $\Sigma_{\text{true}}\succ 0$ is unknown and must be estimated jointly with $x_{\text{true}}$ based on the collected measurements $\{z_{i}\}_{i=1}^{k}$. For convenience, we formulate the problem in the information form and estimate the noise information (or precision) matrix $P_{\text{true}}\triangleq\Sigma_{\text{true}}^{-1}$. Without loss of generality, we assume a non-informative prior on $x_{\text{true}}$ which is almost always the case in real-world applications (effectively treating it as an unknown “parameter”). We assume a Wishart prior on the noise information matrix $P_{\text{true}}$ and denote its probability density function with $\mathcal{W}(P;V,\nu)$ where $V\in\mathbb{S}^{m}_{\succ 0}$ is the scale matrix and the integer $\nu\geq m+1$ is the number of degrees of freedom; Figure 1 illustrates random samples drawn from $\mathcal{W}(P;V,\nu)$. The Wishart distribution is the standard choice for prior in Bayesian statistics for estimating the information matrix from multivariate Gaussian data (in part due to conjugacy); see, e.g., [6, Eq. (2.155)]. In Algorithm 1, we propose a procedure for setting the parameters of the prior, $V$ and $\nu$, when the user has access to a prior estimate $\Sigma_{0}$ for the noise covariance matrix $\Sigma_{\text{true}}$ (e.g., from prior calibration by the manufacturer of the sensor); see Appendix A for a justification.
@@
[…]
@@
$\begin{aligned} &\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}&&\,\,F(x,P)\\ &\text{subject to}&&\text{$P$ is diagonal}.\end{aligned}$ (6)
In the second variant, we constrain the eigenvalues of $\Sigma=P^{-1}$ to $\lambda_{\text{min}},\lambda_{\text{max}}]$ where $\lambda_{\text{max}}\geq\lambda_{\text{min}}>0$. These eigenvalues specify the minimum and maximum variance along all directions in $\mathbb{R}^{m}$ (i.e., variance of all normalized linear combinations of noise components), Therefore, this constraint allows the user to incorporate prior knowledge about the sensor’s noise limits. We will also show that in many real-world instances (especially with a weak or no prior), constraining the smallest eigenvalue of $\Sigma$ (or largest eigenvalue of $P$) is essential for preventing $\Sigma$ from collapsing to zero.
@@
$\begin{aligned} &\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}&&\,\,F(x,P)\\ &\text{subject to}&&\lambda_{\text{max}}^{-1}I\preceq P\preceq\lambda_{\text{min}}^{-1}I.\end{aligned}$ (7)
@@
$\begin{aligned} &\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}&&\,\,F(x,P)\\ &\text{subject to}&&\lambda_{\text{max}}^{-1}I\preceq P\preceq\lambda_{\text{min}}^{-1}I,\\ &&&\text{$P$ is diagonal}.\end{aligned}$ (8)
Remark 1 (Joint ML Estimation).
One can resort to joint ML estimation when a prior distribution is not available. The negative log-likelihood cost function arising in the joint ML estimation problem and its constrained variants (i.e., with eigenvalue and/or diagonal constraints) takes a form similar to $F(x,P)$ defined in (4):
$F_{\text{ML}}(x,P)\triangleq-\log\det P+\langle S(x),P\rangle,$ (9)
where $S(x)$ denotes the sample covariance matrix (3). The unconstrained joint ML problem was also derived in [33].
Remark 2 (Fixing Noise Covariance).
It is worth noting that by fixing $P=P_{0}$, Problem 1 reduces to nonlinear least squares over $x$ and can be (locally) solved using existing solvers [3, 13, 20]. To see this, note that the only term in the objective function (4) that is a function of $x$ can be written as:
$\langle M(x),P_{0}\rangle$ $={\gamma}^{-1}\langle kS(x),P_{0}\rangle+\text{const}$ (10)
$={\gamma}^{-1}\left\langle\sum_{i=1}^{k}r_{i}(x)r_{i}(x)^{\top},P_{0}\right\rangle+\text{const}$ (11)
$=\gamma^{-1}\sum_{i=1}^{k}\|r_{i}(x)\|^{2}_{P_{0}}+\text{const},$ (12)
where $\gamma\triangleq k+\nu-m-1$ is the constant that appears in the denominator of (5).